# Practical Dirac Majorana confusion theorem: Issues and Applicability

### v1

@@ Abstract
quantum statistical property of Majorana neutrinos, practical Dirac Majorana confusion theorem, model-independent formalism of pDMCT, back-to-back muon special kinematic configuration
@@ 1 Introduction
Are neutrinos distinct from their antiparticles like the rest of the known fermions of the Standard Model (SM), or are the neutrino and antineutrino quantum mechanically identical to one another? An affirmative response to the first (second) question would imply neutrinos are Dirac (Majorana) fermions. We are yet to have a definite answer to this fundamental question regarding Dirac or Majorana nature of neutrinos. However, the literature is replete with attempts made on both theoretical and experimental fronts, without much success. The situation is such that there is an ossified belief in the community that the difference between Dirac and Majorana neutrinos via any kinematical observable would be practically impossible to determine due to fact that the observable difference between Dirac and Majorana neutrinos is proportional to the tiny neutrino mass. This is often cited as the “practical Dirac Majorana confusion theorem” (pDMCT, in short) Kayser:1982br  [^1] A few notable precursors to the formulation of pDMCT were the analysis made by Ref. Schechter:1980gk ; Schechter:1981hw . While the “theorem” has been verified in some cases, there is no general model-independent and process-independent proof. Also there is a general lack of clarity regarding its domain of validity. It is therefore necessary and important to explore whether there are any SM allowed processes and kinematic observables that can directly probe the Majorana nature of neutrinos avoiding this pDMCT constraint. In this article we discuss the domain of validity of pDMCT as well as its exceptions. This does not invalidate pDMCT but brings more clarity with regard to its applicability as an useful tool.
@@
For the first time in Refs. Kim:2022xjg and Kim:2021dyj , we respectively implemented both model-independent and specific process-independent studies of pDMCT. The current work borrows some of the features from both the above papers in a more accessible manner and directly addresses some pedagogical aspects to bring clarity. Thus in Sec. 2 we first consider a generic model-independent analysis of processes that contain a neutrino antineutrino pair in the final state. This structure in the final state allows for application of Pauli principle through anti-symmetrisation in the case of Majorana neutrinos. The application of the exclusion principle is independent of the size of the non-zero mass of the neutrino or any other dimensional parameter for that matter. We refer to this as model-independent in the sense that our analysis includes both the SM and new physics (NP) contributions. However, the process itself is allowed in the SM. This is exclusively discussed in Sec. 3. In Sec. 4 we highlight the domain of applicability of pDMCT. This is followed by some pedagogical explanations on pDMCT in Sec. 5. Finally we conclude highlighting the important features in Sec. 6.
@@ 2.1 Details on the process under consideration
It should be noted that in this work we discuss processes where the effect of measurements does not destroy the identical nature of Majorana neutrino and antineutrino. This is akin to putting the constraint that in a double-slit experiment, meant to observe the interference of light, no measurement should identify the slit through which the photon has passed.
@@
$\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}=\frac{1}{2}\Bigg{(}\underbrace{\left|\mathscr{M}(p_{1},p_{2})\right|^{2}}_{\textrm{Directterm}}-\underbrace{\left|\mathscr{M}(p_{2},p_{1})\right|^{2}}_{\textrm{Exchange term}}\Bigg{)}+\underbrace{\textrm{Re}\left(\mathscr{M}(p_{1},p_{2})^{*}\,\mathscr{M}(p_{2},p_{1})\right)}_{\textrm{Interference term}}.$ (3)
@@
$\iint\left(\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}\right)\mathrm{d}^{4}p_{1}\,\mathrm{d}^{4}p_{2}=\iint\underbrace{\textrm{Re}\left(\mathscr{M}(p_{1},p_{2})^{*}\,\mathscr{M}(p_{2},p_{1})\right)}_{\textrm{Interference term}}\mathrm{d}^{4}p_{1}\,\mathrm{d}^{4}p_{2},$ (7)
@@
as $p_{1}$ and $p_{2}$ act as dummy variables since the range of integration is identical.
In the simple processes with the SM mediated interaction alone (e.g. weak neutral current mediated decay $Z^{(*)}\to\nu\bar{\nu}$) one finds that, (1) the “Direct term” and “Exchange term” are equal (see section 2.3), (2) the “Interference term” is proportional to $m_{\nu}^{2}$ and (3) the observable usually requires full phase space integration over $p_{1,2}$. This leads to the conclusion that all kinematical observable differences between Dirac and Majorana cases would be proportional to $m_{\nu}^{2}$ This is essentially the statement of the “practical Dirac-Majorana Confusion Theorem” (pDMCT).
Note that our model-independent and process-independent analysis suggests that if (1) the doubly weak charged current processes [^6] For the sequential weak charged current mediated decays that produce neutrino and antineutrino of different flavors, e.g. the radiative leptonic decay $\ell^{-}\to\nu_{\ell}\;\overline{\nu}_{\ell^{\prime}}\;\ell^{\prime-}\;\gamma$, where ,$\ell,\ell^{\prime}\in\{e,\mu,\tau\}$ and one can never have $\ell=\ell^{\prime}$, the $\nu_{\ell}$ and $\overline{\nu}_{\ell^{\prime}}$ can never be considered as identical fermions even if they might indeed be Majorana fermions. We do not consider such processes in our analysis. that possibly lead to a non-zero difference between “Direct term” and “Exchange term” of Eq. (3) and (2) some kinematic configurations could be identified where individual information about $\nu$, $\overline{\nu}$ can be accessed so as to avoid doing the full phase space integration in Eq. (7), then one might avoid pDMCT constraint.
@@ 3 New Physics scenarios and pDMCT in $\boldsymbol{Z^{(*)}\to\nu_{\ell}\,\overline{\nu}_{\ell}}$
$\left|\mathscr{M}^{D}\right|^{2}$ $=\frac{g_{Z}^{2}}{3}\Bigg{(}\left((C_{V}^{\ell})^{2}+(C_{A}^{\ell})^{2}\right)\left(m_{Z}^{2}-m_{\nu}^{2}\right)+3\left((C_{V}^{\ell})^{2}-(C_{A}^{\ell})^{2}\right)m_{\nu}^{2}\Bigg{)},$ (14)
@@
$\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}$ $=\frac{g_{Z}^{2}}{3}\bigg{(}\left((C_{V}^{\ell})^{2}-(C_{A}^{\ell})^{2}\right)\left(m_{Z}^{2}+2\,m_{\nu}^{2}\right)+6\,(C_{A}^{\ell})^{2}\,m_{\nu}^{2}\bigg{)}$
@@
Note that for the neutral current mediated processes in the SM, such as $Z\to\nu\bar{\nu},~{}e^{+}e^{-}\to\nu\bar{\nu},~{}B\to K\nu\bar{\nu}$ etc., the deduction of energy and/or momentum of the invisible neutrinos does not help to distinguish between Dirac and Majorana neutrinos because the pDMCT holds true even at the amplitude-squared level. In such a case every conceivable observable gives the same result for Dirac and Majorana neutrino, except showing the tiny difference coming from the interference term which is proportional to the neutrino mass. Therefore, only in the presence of physics beyond the SM a distinction may be made between Dirac and Majorana nature of neutrinos through such processesRosen:1982pj ; Kim:2022xjg .
@@ 4.1 The general strategy to probe nature of neutrinos and pDMCT
The general formalism discussed in previous sections may be suitably illustrated by the chart shown in Fig. 1. The red arrows drawn in the figure present the previous works of the pDMCT, all of which have studied the weak neutral current processes within the SM, such as $\gamma^{*}\to\nu\,\bar{\nu}$ Kayser:1982br , $Z\to\nu\,\bar{\nu}$ Shrock:1982jh , $e^{+}e^{-}\to\nu\,\bar{\nu}$ Ma:1989jpa , $K^{+}\to\pi^{+}\,\nu\,\bar{\nu}$ Nieves:1985ir , $e^{+}e^{-}\to\nu\,\bar{\nu}\,\gamma$ Chhabra:1992be , $\left|es\right\rangle\to\left|gs\right\rangle+\nu\,\bar{\nu}\,\gamma$ Yoshimura:2006nd , $e^{-}\gamma\to e^{-}\,\nu\,\bar{\nu}$ Berryman:2018qxn etc. As can be easily seen, all of these processes have confirmed the pDMCT because (1) $\left|\mathscr{M}(p_{1},p_{2})\right|^{2}=\left|\mathscr{M}(p_{2},p_{1})\right|^{2}$ due to the weak neutral current processes within the SM and (2) there is no way to observe or deduce the 4-momenta of $\nu$ and/or $\bar{\nu}$ as these are simple 2- or 3-body processes.
Although there is no general, model-independent, process-independent and observable-independent proof of the “practical Dirac Majorana confusion theorem” [^9] We use the phrase ‘observable-independent proof’ to underline the fact that no mathematical proof of pDMCT can be given without referring to any specific observable. So observables play a pivotal role in the discussion on pDMCT. As there is no observable-independent proof of pDMCT, one is free to explore different possible observables to distinguish between Dirac and Majorana nature of neutrinos in the specific context of one’s chosen process., it is generally assumed to apply to all the probes of Majorana nature of sub-eV neutrinos. The formalism presented in sections 2 and 3 provides a simple model-independent, process-independent and observable independent view of the pathways by which the confusion theorem can be overcome, such as by using the properties of the NP interactions or analysing the “special kinematical scenarios” utilising the chosen parts of neutrino momentum spectra.
@@
4.2 How to overcome pDMCT within the SM by using the special kinematics
in $B^{0}~{}(D,~{}H,~{}J/\psi,~{}\Upsilon(1s),..)\to\mu^{+}\,\mu^{-}\,\nu_{\mu}\,\overline{\nu}_{\mu}$
In most of the experimental scenarios, especially true for processes of the form $X\to Y\,\nu\,\overline{\nu}$, information about individual neutrino momenta is not available. In such a case the difference between the Dirac and the Majorana neutrinos, that may be realised, is given by the integrated interference term in Eq. (7). In such a case the evaluation of the squared Feynman diagram for the “Interference term” in the SM necessarily involves two helicity flips which would make it proportional to $m_{\nu}^{2}$. Thus, if only the SM interactions are considered and one fully integrates over the neutrino and antineutrino 4-momenta, the difference between Dirac and Majorana cases is proportional to $m_{\nu}^{2}$. This may be considered as the most general statement of the “practical Dirac Majorana confusion theorem”.
However, there is no reason a priori for the pDMCT to hold, if one can consider special kinematic configurations where the 4-momenta (or some components of the 4-momenta) of the neutrino and antineutrino are known, so that full integration over 4-momenta is not necessary before comparison with the experiment. This is valid even in the SM. In the following sub-subsections we discuss this scenario where such situation may occur Kim:2021dyj .
@@ 4.2 How to overcome pDMCT within the SM by using the special kinematics
unlike the weak neutral current processes. As explained in section 2.2 if we can measure/deduce the individual energy/momentum of the missing neutrino, then the pDMCT constraint will not apply.
In Fig. 2, we show the general kinematics of $B^{0}\to\mu^{-}\mu^{+}\nu_{\mu}\bar{\nu}_{\mu}$ in the rest frame of $B$. The angles $\theta_{n}$ and $\phi$ are indeed inaccessible in general, as the neutrino pair goes missing. Therefore, for a physically useful differential decay rate we must integrate over both $\theta_{n}$ and $\phi$ i.e.
$\frac{\mathrm{d}^{3}\Gamma^{D/M}}{\mathrm{d}m_{\mu\mu}^{2}\;\mathrm{d}m_{\nu\nu}^{2}\;\mathrm{d}\cos\theta_{m}}$ $=\frac{Y\,Y_{m}\,Y_{n}}{(4\,\pi)^{6}m_{B}^{2}\,m_{\mu\mu}\,m_{\nu\nu}}\int_{-1}^{1}\int_{0}^{2\pi}\Big{\langle}\big{|}\mathscr{M}^{D/M}\big{|}^{2}\Big{\rangle}\,\mathrm{d}\cos\theta_{n}\;\mathrm{d}\phi,$ (17)
where $Y$ is the magnitude of 3-momentum of the di-muon system (muon pair with invariant mass $m_{\mu\mu}$) or di-neutrino system (neutrino pair with invariant mass $m_{\nu\nu}$) in the rest frame of $B^{0}$, $Y_{m}$ is the magnitude of 3-momentum of $\mu^{\pm}$ in the di-muon rest frame, $Y_{n}$ is the magnitude of 3-momentum of $\nu_{\mu}$ or $\overline{\nu}_{\mu}$ in the di-neutrino rest frame. It is straightforward to show that the difference between Dirac and Majorana cases is given by
@@
where $m_{B}$ is the mass of the $B^{0}$ meson. Let us choose $E_{\mu}$ as one independent variable. The other independent variable would then be the angle, say $\theta$, between the muon direction and the neutrino direction.
@@
Whenever $E_{1}=E_{2}$ for any value of the angle $\Theta$ between the neutrino and antineutrino we get $\cos\theta_{n}=0$. By analytic continuation we extend this feature to the back-to-back kinematics for which the $\cos\theta_{n}$ has a discontinuity otherwise. Moreover, in the back-to-back case we have both the back-to-back muons and the back-to-back neutrino antineutrino pair, in one single plane. This implies that for the back-to-back case we have $\phi=0$. These choices put the orientation of the coordinate axes in such a way that the back-to-back neutrino and antineutrino fly away defining the $x$-axis. The $xz$-plane in Fig. 2 is the one in which the $3$-momenta of muons lie, and now the back-to-back neutrino antineutrino define the $x$-direction. The direction perpendicular to the neutrino direction is the $z$-direction. If we define the angle between the neutrino and muon directions to be $\theta$, then $\theta_{m}=\pi/2-\theta$. This implies that

### v4

@@ Abstract
quantum statistical property of Majorana neutrinos, practical Dirac Majorana confusion theorem, model-independent formalism of pDMCT, back-to-back muon special kinematic configuration, non-existence of massless chiral Majorana fermion
journal: Eur. Phys. J. C (2023) 83:972
@@ 1 Introduction
Are neutrinos distinct from their antiparticles like the rest of the known fermions of the Standard Model (SM), or are the neutrino and antineutrino quantum mechanically identical to one another? An affirmative response to the first (second) question would imply neutrinos are Dirac (Majorana) fermions. We are yet to have a definite answer to this fundamental question regarding Dirac or Majorana nature of neutrinos. However, the literature is replete with attempts made on both theoretical and experimental fronts, without much success. The situation is such that there is an ossified belief in the community that the difference between Dirac and Majorana neutrinos via any kinematical observable would be practically impossible to determine due to fact that the observable difference between Dirac and Majorana neutrinos is proportional to the tiny neutrino mass. This is often cited as the “practical Dirac Majorana confusion theorem” (pDMCT, in short) Kayser:1982br  [^1] A few notable precursors to the formulation of pDMCT were the analysis made by Refs. Schechter:1980gk ; Schechter:1981hw .. While the “theorem” has been verified in some cases, there is no general model-independent and process-independent proof. Also there is a general lack of clarity regarding its domain of validity. It is therefore necessary and important to explore whether there are any SM allowed processes and kinematic observables that can directly probe the Majorana nature of neutrinos avoiding this pDMCT constraint. In this article we discuss the domain of validity of pDMCT as well as its exceptions. This does not invalidate pDMCT but brings more clarity with regard to its applicability as an useful tool.
@@
For the first time in Refs. Kim:2022xjg and Kim:2021dyj , we respectively implemented both model-independent and specific process-independent studies of pDMCT. The current work borrows some of the features from both the above papers in a more accessible manner and directly addresses some pedagogical aspects to bring clarity. Thus in Sec. 2 we first consider a generic model-independent analysis of processes that contain a neutrino antineutrino pair in the final state. This structure in the final state allows for application of Pauli exclusion principle through anti-symmetrisation in the case of Majorana neutrinos. The application of the exclusion principle is independent of the size of the non-zero mass of the neutrino or any other dimensional parameter for that matter. We refer to this as model-independent in the sense that our analysis includes both the SM and new physics (NP) contributions. However, the process itself is allowed in the SM. This is exclusively discussed in Sec. 3. In Sec. 4 we highlight the domain of applicability of pDMCT. This is followed by some pedagogical explanations on pDMCT in Sec. 5. Finally we conclude highlighting the important features in Sec. 6.
@@ 2.1 Details on the process under consideration
It should be noted that in this work we discuss processes where the effect of measurements does not destroy the identical nature of Majorana neutrino and antineutrino. This is akin to putting the constraint that in a double-slit experiment, meant to observe the interference of light, no measurement should identify the slit through which the photon has passed. We again stress that direct observation of the state of neutrinos (via neutrino detector) destroys the effect of quantum statistics since the neutrinos after the detection are projected on to specific helicity states (and thus no longer remain indistinguishable) by observation. As is well known, quantum statistics requires absolutely identical indistinguishable particles.
@@
$\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}$ $=\frac{1}{2}\Bigg{(}\underbrace{\left|\mathscr{M}(p_{1},p_{2})\right|^{2}}_{\textrm{Direct term}}-\underbrace{\left|\mathscr{M}(p_{2},p_{1})\right|^{2}}_{\textrm{Exchange term}}\Bigg{)}$
$\quad+\underbrace{\textrm{Re}\left(\mathscr{M}(p_{1},p_{2})^{*}\,\mathscr{M}(p_{2},p_{1})\right)}_{\textrm{Interference term}}.$ (3)
@@
$\iint\left(\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}\right)\mathrm{d}^{4}p_{1}\,\mathrm{d}^{4}p_{2}$
$=\iint\underbrace{\textrm{Re}\left(\mathscr{M}(p_{1},p_{2})^{*}\,\mathscr{M}(p_{2},p_{1})\right)}_{\textrm{Interference term}}\mathrm{d}^{4}p_{1}\,\mathrm{d}^{4}p_{2},$ (7)
@@
as $p_{1}$ and $p_{2}$ act as dummy variables since the range of integration is identical. Therefore, if we are able to find any smart method to deduce (or infer) energy-momenta of missing neutrinos, in order not to make full phase space integral over missing neutrinos, and if there exists non-trivial difference between the “Direct term" and the “Exchange term" as shown in Eq. (4), the pDMCT would be nicely avoided.
In the simple processes with the SM mediated interaction alone (e.g. weak neutral current mediated decay $Z^{(*)}\to\nu\bar{\nu}$) one finds that, (1) the “Direct term” and “Exchange term” are equal (see section 2.3), (2) the “Interference term” is proportional to $m_{\nu}^{2}$ and (3) the observable usually requires full phase space integration over $p_{1,2}$. This leads to the conclusion that all kinematical observable differences between Dirac and Majorana cases would be proportional to $m_{\nu}^{2}$. This is essentially the statement of the “practical Dirac-Majorana Confusion Theorem” (pDMCT).
Note that our model-independent and process-independent analysis suggests that if (1) the doubly weak charged current processes [^6] For the sequential weak charged current mediated decays that produce neutrino and antineutrino of different flavors, e.g. $\ell^{-}\to\nu_{\ell}\;\overline{\nu}_{\ell^{\prime}}\;\ell^{\prime-}$, where ,$\ell,\ell^{\prime}\in\{e,\mu,\tau\}$ and one can never have $\ell=\ell^{\prime}$, the $\nu_{\ell}$ and $\overline{\nu}_{\ell^{\prime}}$ can never be considered as identical fermions even if they might indeed be Majorana fermions. We do not consider such processes in our analysis. that possibly lead to a non-zero difference between “Direct term” and “Exchange term” of Eq. (3) and (2) some kinematic configurations could be identified where individual information about $\nu$, $\overline{\nu}$ can be accessed so as to avoid doing the full phase space integration in Eq. (7), then one might avoid pDMCT constraint.
@@ 3 New Physics scenarios and pDMCT in $\boldsymbol{Z^{(*)}\to\nu_{\ell}\,\overline{\nu}_{\ell}}$
$\left|\mathscr{M}^{D}\right|^{2}$ $=\frac{g_{Z}^{2}}{3}\Bigg{(}\left((C_{V}^{\ell})^{2}+(C_{A}^{\ell})^{2}\right)\left(m_{Z}^{2}-m_{\nu}^{2}\right)$
$\hskip 56.9055pt+3\left((C_{V}^{\ell})^{2}-(C_{A}^{\ell})^{2}\right)m_{\nu}^{2}\Bigg{)},$ (14)
@@
$\left|\mathscr{M}^{D}\right|^{2}-\left|\mathscr{M}^{M}\right|^{2}$ $=\frac{g_{Z}^{2}}{3}\bigg{(}\left((C_{V}^{\ell})^{2}-(C_{A}^{\ell})^{2}\right)\left(m_{Z}^{2}+2\,m_{\nu}^{2}\right)$
$\hskip 28.45274pt+6\,(C_{A}^{\ell})^{2}\,m_{\nu}^{2}\bigg{)}$
@@
Note that for the neutral current mediated processes in the SM, such as $Z\to\nu\bar{\nu},~{}e^{+}e^{-}\to\nu\bar{\nu},~{}B\to K\nu\bar{\nu}$ etc., the deduction of energy and/or momentum of the invisible neutrinos does not help to distinguish between Dirac and Majorana neutrinos because the pDMCT holds true even at the amplitude-squared level. In such a case every conceivable observable gives the same result for Dirac and Majorana neutrino, except showing the tiny difference coming from the interference term which is proportional to the neutrino mass. Therefore, only in the presence of physics beyond the SM Rosen:1982pj ; Rodejohann:2017vup ; Kim:2022xjg a distinction may be made between Dirac and Majorana nature of neutrinos through such processes.
@@ 4.1 The general strategy to probe nature of neutrinos and pDMCT
The general formalism discussed in previous sections may be suitably illustrated by the chart shown in Fig. 1. The red (color in online) arrows drawn in the figure present the previous works of the pDMCT, all of which have studied the weak neutral current processes within the SM, such as $\gamma^{*}\to\nu\,\bar{\nu}$ Kayser:1982br , $Z\to\nu\,\bar{\nu}$ Shrock:1982jh , $e^{+}e^{-}\to\nu\,\bar{\nu}$ Ma:1989jpa , $K^{+}\to\pi^{+}\,\nu\,\bar{\nu}$ Nieves:1985ir , $e^{+}e^{-}\to\nu\,\bar{\nu}\,\gamma$ Chhabra:1992be , $\left|es\right\rangle\to\left|gs\right\rangle+\nu\,\bar{\nu}\,\gamma$ Yoshimura:2006nd , $e^{-}\gamma\to e^{-}\,\nu\,\bar{\nu}$ Berryman:2018qxn , $\nu+N\to N\,\nu\,\gamma$ Millar:2018hkv , etc. As can be easily seen, all of these processes have confirmed the pDMCT because (1) $\left|\mathscr{M}(p_{1},p_{2})\right|^{2}=\left|\mathscr{M}(p_{2},p_{1})\right|^{2}$ due to the weak neutral current processes within the SM and/or (2) there is no way to observe or deduce the 4-momenta of $\nu$ and/or $\bar{\nu}$ as these are simple 2- or 3-body processes.
Although there is no general, model-independent, process-independent and observable-independent proof [^9] We use the phrase ‘observable-independent proof’ to underline the fact that no mathematical proof of pDMCT can be given without referring to any specific observable. So observables play a pivotal role in the discussion on pDMCT. As there is no observable-independent proof of pDMCT, one is free to explore different possible observables to distinguish between Dirac and Majorana nature of neutrinos in the specific context of one’s chosen process. of the “practical Dirac Majorana confusion theorem”, it is generally assumed to apply to all the probes of Majorana nature of sub-eV neutrinos. The formalism presented in sections 2 and 3 provides a simple model-independent, process-independent and observable independent view of the pathways by which the confusion theorem can be overcome, such as by using the properties of the NP interactions or analysing the “special kinematical scenarios” utilising the chosen parts of neutrino momentum spectra.
@@
4.2 How to overcome pDMCT within the SM by using the special kinematics in $\boldsymbol{B^{0}~{}(D,~{}H,~{}J/\psi,~{}\Upsilon(1s),..)\to\mu^{+}\,\mu^{-}\,\nu_{\mu}\,\overline{\nu}_{\mu}}$
In most of the experimental scenarios, especially true for processes of the form $X\to Y\,\nu\,\overline{\nu}$, information about individual neutrino momenta is not available. In such a case the difference between the Dirac and the Majorana neutrinos, that may be realised, is given by the integrated interference term in Eq. (7). In such a case the evaluation of the squared Feynman diagram for the “Interference term” in the SM necessarily involves two helicity flips which would make it proportional to $m_{\nu}^{2}$. Thus, if only the SM interactions are considered and one fully integrates over the neutrino and antineutrino 4-momenta due to unobservable neutrinos , the difference between Dirac and Majorana cases is proportional to $m_{\nu}^{2}$. This may be considered as the most general statement of the “practical Dirac Majorana confusion theorem”.
However, there is no reason a priori for the pDMCT to hold, if one can consider special kinematic configurations where the 4-momenta (or some components of the 4-momenta) of the neutrino and antineutrino are known, so that full integration over 4-momenta is not necessary before comparison with the experiment. This is valid even in the SM. In the following sub-subsections we discuss this scenario where such situation may occur.
@@
unlike the weak neutral current processes. As explained in section 2.2 if we can measure or deduce the individual energy or momentum of the missing neutrino, then the pDMCT constraint will not apply.
In Fig. 2, we show the general kinematics of $B^{0}\to\mu^{-}\mu^{+}\nu_{\mu}\bar{\nu}_{\mu}$ in the rest frame of $B$. The angles $\theta_{n}$ and $\phi$ are indeed inaccessible in general, as the neutrino pair goes missing. Therefore, for a physically useful differential decay rate we must integrate over both $\theta_{n}$ and $\phi$, i.e.
$\frac{\mathrm{d}^{3}\Gamma^{D/M}}{\mathrm{d}m_{\mu\mu}^{2}\;\mathrm{d}m_{\nu\nu}^{2}\;\mathrm{d}\cos\theta_{m}}$ $=\frac{Y\,Y_{m}\,Y_{n}}{(4\,\pi)^{6}m_{B}^{2}\,m_{\mu\mu}\,m_{\nu\nu}}$
×$\quad\times\int_{-1}^{1}\int_{0}^{2\pi}\Big{\langle}\big{|}\mathscr{M}^{D/M}\big{|}^{2}\Big{\rangle}\,\mathrm{d}\cos\theta_{n}\;\mathrm{d}\phi,$ (17)
where $Y$ is the magnitude of 3-momentum of the di-muon system (muon pair with invariant mass $m_{\mu\mu}$) or di-neutrino system (neutrino pair with invariant mass $m_{\nu\nu}$) in the rest frame of $B^{0}$, $Y_{m}$ is the magnitude of 3-momentum of $\mu^{\pm}$ in the di-muon rest frame, $Y_{n}$ is the magnitude of 3-momentum of $\nu_{\mu}$ or $\overline{\nu}_{\mu}$ in the di-neutrino rest frame. It is straightforward to show that the difference between Dirac and Majorana cases, as shown in Eqs. (33, 34) of Kim:2021dyj , is given by
@@
where $m_{B}$ is the mass of the $B^{0}$ meson. Let us choose $E_{\mu}$ as one independent variable. The other independent variable would then be the angle, say $\theta$, between the muon direction and the neutrino direction. We would like again to point out that direct observation of neutrinos destroys the effect of quantum statistics, as explained before. Hence, any inference from qauntum statistical probe necessarily requires no direct detection of the final neutrinos Kim:2024xof . This is precisely achieved in the back-to-back special kinematic configuration where the energy momentum information is inferred purely by observing the muons.
@@
Whenever $E_{1}=E_{2}$ for any value of the angle $\Theta$ between the neutrino and antineutrino we get $\cos\theta_{n}=0$. By analytic continuation we extend this feature to the back-to-back kinematics for which the $\cos\theta_{n}$ has a discontinuity otherwise. Moreover, in the back-to-back case we have both the back-to-back muons and the back-to-back neutrino antineutrino pair, in one single plane. This implies that for the back-to-back case we have $\phi=0$ Kim:2023ohr . These choices put the orientation of the coordinate axes in such a way that the back-to-back neutrino and antineutrino fly away defining the $x$-axis. The $xz$-plane in Fig. 2 is the one in which the $3$-momenta of muons lie, and now the back-to-back neutrino antineutrino define the $x$-direction. The direction perpendicular to the neutrino direction is the $z$-direction.
The back-to-back kinematics specified by $\vec{p}_{+}+\vec{p}_{-}=\vec{0}=\vec{p}_{1}+\vec{p}_{2}$ has $3$ additional constraint equations, which implies that only $2$ out of the initial $5$ variables would remain independent in back-to-back case, i.e. $3$ variables should either be fixed by the back-to-back conditions or they must be dependent on the $2$ variables which are independent (ie. $E_{\mu},~{}\cos\theta_{m}$) Kim:2023ohr . If we define the angle between the neutrino and muon directions to be $\theta$, then $\theta_{m}=\pi/2-\theta$. This implies that